OpenStudy (anonymous):

Find x and y such that : xy+(x+y)=19 and xy(x+y)=84.

OpenStudy (jamesj):

Let a = xy and b = x+y. Then a + b = 19 ab = 84. Write b = 19-a. Then substituting into the second equation, a(19-a) = 84 => a^2 - 19a + 84 = 0 => a = 12 or 7 For a = 12, b = 7; and for a = 7, b = 12. We expected that btw, because the equations in a and b are symmetric. Now look at the equations for x and y, given a = 12, b = 7. xy = a = 12 x+y = b = 7 => y = 7 - x. Substitute that equation for y now in the equation xy = 12. Then solve.